OpenStudy (anonymous):

Solve for x. 3√2x / 4= 12

OpenStudy (anonymous):

3x√2 = 48 x = 48/(3√2)

OpenStudy (lgbasallote):

\(\huge \frac{3\sqrt{2x}}{4} = 12\) \(\huge 3\sqrt{2x} = 48\) \(\huge \sqrt{2x} = \frac{48}{3}\) \(\huge \sqrt{2x} = 16\) \(\huge 2x = 256\) \(\huge x = 128\)

OpenStudy (anonymous):

I like the way you typed that haha and thanks :)

OpenStudy (lgbasallote):

\(\Huge \color{maroon}{\mathbb{\text{haha thanks <tips hat>}}}\)

OpenStudy (anonymous):

:D